\begin{thebibliography}{00}
\bibitem{Amri-Rachdi-1} Amri, B., Rachdi, L.T.: Beckner logarithmic uncertainty principle for the Riemann-Liouville operator. International Journal of Mathematics, 24(09), 1350070. doi:10.1142/s0129167x13500705 (2013)

\bibitem{Amri-Rachdi-2} Amri, B., Rachdi, L.T.: Uncertainty Principle in Terms of Entropy for the Riemann–Liouville Operator. Bulletin of the Malaysian Mathematical Sciences Society, 39(1), 457–481. doi:10.1007/s40840-015-0121-5 (2015)

\bibitem{Arfaoui-Rezgui-BenMabrouk} S. Arfaoui, I. Rezgui and A. Ben Mabrouk, Wavelet Analysis on the Sphere: Spheroidal Wavelets. Walter de Gruyter (March 20, 2017), ISBN-10: 311048109X, ISBN-13: 978-3110481099.

\bibitem{Arfaoui-BenMabrouk1} Arfaoui, S., Ben Mabrouk, A. Some Ultraspheroidal Monogenic Clifford Gegenbauer Jacobi Polynomials and Associated Wavelets. Adv. Appl. Clifford Algebras 27, 2287–2306 (2017). https://doi.org/10.1007/s00006-017-0788-9.

\bibitem{Arfaoui-BenMabrouk2} Arfaoui, S., Ben Mabrouk, A. Some Old Orthogonal Polynomials Revisited and Associated Wavelets: Two-Parameters Clifford-Jacobi Polynomials and Associated Spheroidal Wavelets. Acta Appl Math 155, 177–195 (2018). https://doi.org/10.1007/s10440-017-0150-1.

\bibitem{Arfaoui-BenMabrouk-Cattani1} Arfaoui, S., Ben Mabrouk, A. Cattani, C. New Type of Gegenbauer–Hermite Monogenic Polynomials and Associated Clifford Wavelets. J Math Imaging Vis 62, 73–97 (2020). https://doi.org/10.1007/s10851-019-00914-y.

\bibitem{Arfaoui-BenMabrouk-Cattani2} Arfaoui, S., Ben Mabrouk, A. Cattani, C. New Type of Gegenbauer-Jacobi-Hermite Monogenic Polynomials and Associated Continuous Clifford Wavelet Transform. Acta Appl Math (2020). https://doi.org/10.1007/s10440-020-00322-0

\bibitem{Banouh} Banouh, H., Ben Mabrouk, A. and Kesri, M., Clifford Wavelet Transform and the Uncertainty Principle. Adv. Appl. Clifford Algebras 29, 106 {2019}, 23 pqges. https://doi.org/10.1007/s00006-019-1026-4.

\bibitem{BrackxChisholmSoucek} Brackx, F.,  Chisholm J.S.R. and Soucek, V.:Clifford Analysis and Its Applications. NATO Science Series, Series II: Mathematics, Physics and Chemistry - Vol. 25. Springer (2000)

\bibitem{Brackx-Schepper-Sommen0} Brackx, F., Delanghe, R. and Sommen, F.: Clifford analysis, Pitman Publication, (1982)

\bibitem{Brackxetal2006} Brackx, F.,  De Schepper, N. and Sommen, F.: The Two-Dimensional Clifford-Fourier Transform. J. Math. Imaging, 26, 5-18 (2006)

\bibitem{Brackx-Schepper-Sommen2} Brackx, F.,  De Schepper, N. and Sommen, F.: The Fourier Transform in Clifford analysis. Advances in Imaging and Electron Physics, 156, 55-201 (2009)

\bibitem{Brackxetal2006a} Brackx, F.,  De Schepper, N. and Sommen, F.: Clifford-Hermite and two-dimensional Clifford-Gabor filters for early vision. In (digital) Proceedings 17th International Conference on the Application of Computer Science and Mathematics in Architecture and Civil Engineering (K. G\"urlebeck and C. K\"onke, eds.), Bauhaus-Universit\"at Weimar, July 12–14 (2006)

\bibitem{Brackxetal2013} Brackx, F., Hitzer, E., Sangwine, S.J.: History of quaternion and Clifford-fourier transforms and wavelets, in Quaternion and Clifford fourier transforms and wavelets. In Trends in Mathematics 27. p.XI-XXVII (2013)

\bibitem{Brackx2001a} Brackx, F., Sommen, F.: \newblock The continuous wavelet transform in Clifford analysis. \newblock In {\em Clifford Analysis and Its Applications}, pages 9--26. Springer Netherlands, (2001)
	
\bibitem{Bujacketal} Bujack, R., De Bie, H., De Schepper, N., Scheuermann, G.: Convolution Products for Hypercomplex Fourier Transforms. J Math Imaging Vis, 19 pages, DOI 10.1007/s10851-013-0430-y (2013)

\bibitem{Bujacketal1} Bujack, R., Scheuermann, G., Hitzer, E.: A General Geometric Fourier Transform, in E. Hitzer, S.J. Sangwine (eds.), Quaternion and Clifford
Fourier transforms and wavelets, Trends in Mathematics 27, Birkhauser, Basel, pp. 155-176. DOI: 10.1007/978-3-0348-0603-$9_-$8 (2013)

\bibitem{Bujacketal2} Bujack, R., Scheuermann, G., Hitzer, E.: A General Geometric Fourier Transform Convolution Theorem, Adv. Appl. Clifford Alg., Vol. 23(1), 15-38, DOI: 10.1007/s00006-012-0338-4 (2013)

\bibitem{Clifford1882} Clifford, W.K.: \newblock On the classification of geometric algebras. \newblock {\em Mathematical Papers}, pages 397--401, (1882)
	
\bibitem{Dahkleteal} Dahkle, S., Kutyniok, G., Maass, P., Sagiv, C., Stark, H.-G., Teschke, G.: The uncertainty principle associated with the continuous shearlet transform. International Journal of Wavelets, Multiresolution and Information Processing, 6(2), 157–181 (2008)

\bibitem{Delanghe} Delanghe, R.: Clifford Analysis: History and Perspective, Computational Methods and Function Theory, 1(1), 107-153 (2001)

\bibitem{DeBie1} De Bie, H.: Clifford algebras, Fourier transforms and quantum mechanics, arXiv: 1209.6434v1, 39 pages (2012)

\bibitem{DeBie2} De Bie, H., Xu, Y.: On the Clifford-Fourier transform. ArXiv: 1003.0689, 30 pages (2010)

\bibitem{DeSchepper2006} De Schepper, N.: \newblock {\em Multi-dimensional continuous wavelet transforms and generalized Fourier transforms in Clifford analysis}. \newblock PhD thesis, Ghent University (2006)

\bibitem{Dian} Dian Tunjung, N., Zainal Arifin, A., Soelaiman, R.: Medical Image Segmentation Using Generalized Gradient Vector Flow and Clifford Geometric Algebra. International Conference on Biomedical Engineering, Surabaya, Indonesia, November 11, 5 pages (2008)

\bibitem{Dirac1928} Dirac, P.A.M.: \newblock The quantum theory of the electron. \newblock {\em Proceedings of the Royal Society of London. Series A, Containing Papers of a Mathematical and Physical Character}, 117(778), 610--624, (1928).
	
\bibitem{ElHaouietal} El Haoui, Y., Fahlaoui, S., Hitzer, E.: Generalized Uncertainty Principles associated with the Quaternionic Offset Linear Canonical Transform. arXiv:1807.04068v2 [math.CA], 15 pages (2019)

\bibitem{ElHaoui-Fahlaoui} El Haoui, Y., Fahlaoui, S.: Donoho-Stark's Uncertainty Principles in Real Clifford Algebras. arXiv:1902.08465v1 [math.CA], 9 pages (2019)

\bibitem{Feichtinger-Grochenig} Feichtinger, H.G., Gr\"ochenig, K.: Gabor Wavelets and the Heisenberg Group: Gabor Expansions and Short Time Fourier Transform from the Group Theoretical Point of View. Wavelets, 359–397. doi:10.1016/b978-0-12-174590-5.50018-6 (1992)

\bibitem{Grossmann1986} Grossmann, A., Morlet, J., Paul, T.: {\em Transforms associated to square integrable group representations. II : examples}. \newblock Annales de l'IHP Physique th\'eorique, 45(3), 293--309 (1986)

\bibitem{Grossmann1984} Grossmann, A., Morlet, J.: \newblock Decomposition of hardy functions into square integrable wavelets of constant shape. \newblock {\em SIAM Journal on Mathematical Analysis}, 15, 723--736, (1984)

\bibitem{Hamilton1844} Hamilton, W.R.: \newblock On a new species of imaginary quantities connected with a theory of quaternions. \newblock In {\em Proceedings of the Royal Irish Academy} 2, 424--434 (1844)
	
\bibitem{Hamilton1866} Hamilton, W.R.: \newblock {\em Elements of quaternions}. \newblock Longmans, Green, \& Company (1866)
	
\bibitem{Heisenberg1927} Heisenberg, W.: \newblock \"Uber den anschaulichen inhalt der quantentheoretischen kinematik und mechanik. \newblock {\em Zeitschrift fur Physik} 43, 172--198 (1927).
	
\bibitem{Heisenberg1985} W. Heisenberg. \newblock \"Uber den anschaulichen inhalt der quantentheoretischen kinematik und mechanik. \newblock In {\em Original Scientific Papers Wissenschaftliche Originalarbeiten}, 478--504, Springer Berlin Heidelberg (1985)
	
\bibitem{Hitzer} Hitzer, E.: New Developments in Clifford Fourier Transforms, in N. E. Mastorakis, P. M. Pardalos, R. P. Agarwal, L. Kocinac (eds.), Advances in Applied and Pure Mathematics, Proceedings of the 2014 International Conference on Pure
Mathematics, Applied Mathematics, Computational Methods (PMAMCM 2014), Santorini Island, Greece, July 17-21, 2014, Mathematics and Computers in Science and Engineering Series 29, 19-25 (2014)

\bibitem{Hitzer1} Hitzer, E.: Clifford (Geometric) Algebra Wavelet Transform, in V. Skala and D. Hildenbrand (eds.), Proc. of GraVisMa 2009, 02-04 Sep. 2009, Plzen, Czech Republic, pp. 94-101 (2009)

\bibitem{Hitzer2} Hitzer, E.: Directional Uncertainty Principle for Quaternion Fourier Transform. Adv. appl. Clifford alg. 20, 271–284. DOI: 10.1007/s00006-009-0175-2 (2010)

\bibitem{Hitzer-Mawardi-1} Hitzer, E., Mawardi, B.: Uncertainty Principle for the Clifford-Geometric Algebra $\mathcal{Cl}_{3,0}$ based on Clifford Fourier Transform. arXiv:1306.2089v1 [math.RA] 4 pages (2013)

\bibitem{Hitzer-Tachibana} Hitzer, E.: Tutorial on Fourier Transformations and Wavelet Transformations in Clifford Geometric Algebra, in K. Tachibana (ed.), Lecture notes of the International Workshop for Computational Science with Geometric Algebra (FCSGA2007), Nagoya University, Japan, 14-21 Feb. 2007, pp. 65-87 (2007)

\bibitem{Jorgensen} Jorgensen, P., Tian, F.: Non-commutative analysis. World Scientific Publishing Company (2017)

\bibitem{Kouetal} Kou, K.I., Ou, J.-Y., Morais, J.: On Uncertainty Principle for Quaternionic Linear Canonical Transform. Abstract and Applied Analysis, Article ID 725952, 14 pages, http://dx.doi.org/10.1155/2013/725952 (2013)

\bibitem{Ma-Zhao} Ma, G., Zhao, J.: Quaternion Ridgelet Transform and Curvelet Transform. Adv. Appl. Clifford Algebras 28:80, https://doi.org/10.1007/s00006-018-0897-0 (2018)

\bibitem{Mawardi} Mawardi, B.: Construction of Quaternion-Valued Wavelets, Matematika 26(1), 107-114 (2010)

\bibitem{Mawardi-Ryuichi} Mawardi, B., Ryuichi, A.: A Simplified Proof of Uncertainty Principle for Quaternion Linear Canonical Transform. Abstract and Applied Analysis, Article ID 5874930, 11 pages. http://dx.doi.org/10.1155/2016/5874930 (2016)

\bibitem{Mawardi-Ryuichi-1} Mawardi, B., Ashino, R.: Logarithmic uncertainty principle for quaternion linear canonical transform. Proceedings of the 2016 International Conference on Wavelet Analysis and Pattern Recognition, Jeju, South Korea, 10-13 July, 6 pages (2016)

\bibitem{Mawardi-Ryuichi-2} Mawardi, B., Ashino, R.: A Variation on Uncertainty Principle and Logarithmic Uncertainty Principle for Continuous Quaternion Wavelet
Transforms. Abstract and Applied Analysis, Article ID 3795120, 11 pages, https://doi.org/10.1155/2017/3795120 (2017)

\bibitem{Mawardi-Hitzer-1} Mawardi, B., Hitzer, E.: Clifford Algebra $Cl(3,0)$-valued Wavelets and Uncertainty Inequality for Clifford Gabor Wavelet Transformation, Preprints of Meeting of the Japan Society for Industrial and Applied Mathematics, ISSN: 1345-3378, Tsukuba University, 16-18 Sep. 2006, Tsukuba, Japan, pp. 64-65 (2006)

\bibitem{Mawardi-Hitzer-2} Mawardi, B., Hitzer, E.: Clifford algebra $Cl(3,0)$-valued wavelet transformation, Clifford wavelet uncertainty inequality and Clifford Gabor wavelets. International Journal of Wavelets, Multiresolution and Information Processing 5(6), 997--1019 (2007)

\bibitem{Mawardi-Hitzer-3} Mawardi, B., Hitzer, E.: Clifford Fourier Transformation and Uncertainty Principle for the Clifford Geometric Algebra $Cl_{3,0}$. Adv. appl. Clifford alg. 16, 41--61. DOI 10.1007/s00006-006-0003-x (2006)

\bibitem{Mawardi-Hitzer-4} Mawardi, B., Hitzer, E.: Clifford Fourier Transform on Multivector Fields and Uncertainty Principles for Dimensions $n=2(mod4)$ and $n=3(mod4)$. Adv. appl. Clifford alg. 18, 715--736. DOI: 10.1007/s00006-008-0098-3 (2008)

\bibitem{Mawardi-Hitzer-Hayashi-Ashino} Mawardi, B., Hitzer, E., Hayashi, A., Ashino, R.: An Uncertainty Principle for Quaternion Fourier Transform, Computer \& Mathematics with Applications 56, 2398-2410 (2008)

\bibitem{Mawardi-Ashino-Vaillancourt} Mawardi, B., Ashino, R., Vaillancourt, R.: Two-dimensional quaternion wavelet transform. Applied Mathematics and Computation 218, 10--21 (2011)

\bibitem{Nagata} Nagata, K., Nakamura, T: Violation of Heisenberg's Uncertainty
Principle. Open Access Library Journal, 2, e1797. http://dx.doi.org/10.4236/oalib.1101797 (2015)

\bibitem{Rachdi-Meherzi} Rachdi, L.T., Meherzi, F.: Continuous Wavelet Transform and Uncertainty Principle Related to the Spherical Mean Operator. Mediterranean Journal of Mathematics, 14(1). doi:10.1007/s00009-016-0834-1 (2016)

\bibitem{Rachdi-Amri-Hammami} Rachdi, L.T., Amri, B., Hammami, A.: Uncertainty principles and time frequency analysis related to the Riemann–Liouville operator. Annali Dell'Universita' Di Ferrara. doi:10.1007/s11565-018-0311-9 (2018)

\bibitem{Rachdi-Herch} Rachdi, L.T., Herch, H.: Uncertainty principles for continuous wavelet transforms related to the Riemann–Liouville operator. Ricerche Di Matematica, 66(2), 553–578. doi:10.1007/s11587-017-0320-5 (2017)

\bibitem{Rizo} Rizo-Rodríguez, D., Mendez-Vazquez, H., Garcia-Reyes, E.: Illumination Invariant Face Recognition Using Quaternion-Based Correlation Filters. J. Math. Imaging Vis., 45, 164-175 (2013)

\bibitem{Sauetal} Sau, K., Basaka, R.K., Chanda, A.: Image Compression based on Block Truncation Coding using Clifford Algebra. International Conference on Computational Intelligence: Modeling Techniques and Applications (CIMTA), Procedia Technology 10, 699-706 (2013)

\bibitem{Sen} Sen, D.: The uncertainty relations in quantum mechanics. Current Science 107(2), 203--218 (2018)

\bibitem{Sommen2015} Sommen, F, De Schepper, H.: \newblock {\em Introductory Clifford Analysis}, pages 1--27. , Basel (2015)

\bibitem{Soulard} Soulard, R., Carr\'e, P.: Characterization of color images with multiscale monogenic maxima. IEEE Transactions on Pattern Analysis and Machine Intelligence, 40(10), 2289--2302 (2018)

\bibitem{Stabnikov} Stabnikov, P.A.: Geometric Interpretation of the Uncertainty Principle. Natural Science 11(5), 146-148 (2019)

\bibitem{Wietzke} Wietzke, L., Sommer, G.: The Signal Multi-Vector. J. Math. Imaging Vis. 37, 132-150 (2010)

\bibitem{Weyl1950} Weyl, H.: \newblock {\em the theory of groups and quantum mechanics}. \newblock Dover, New York, second edition (1950)

\bibitem{Yangetal1} Yang, Y., Dang, P., Qian, T.: Stronger uncertainty principles
for hypercomplex signals, Complex Variables and Elliptic Equations, 60(12), 1696--1711, DOI: 10.1080/17476933.2015.1041938 (2015)

\bibitem{Yang-Kou} Yang, Y., Kou, K.I.: Uncertainty principles for hypercomplex signals in the linear canonical transformdomains. Signal Processing 95, 67–75 (2014)

\bibitem{Zou} Zou, C., Kou, K. I.: Hypercomplex Signal Energy Concentration in the Spatial and Quaternionic Linear Canonical Frequency Domains, arXiv preprint arXiv:1609.00890 (2016)
\end{thebibliography}

\documentclass{birkjour}
\newtheorem{thm}{Theorem}[section]
\newtheorem{cor}[thm]{Corollary}

\theoremstyle{defn}
\newtheorem{defn}[thm]{Definition}
\theoremstyle{rem}

\numberwithin{equation}{section}

\begin{document}

%
%
%
%
%
%
%
%
%

\title{A sharp Clifford-wavelet Heisenberg-type uncertainty principle}

\author[Banouh]{Hicham Banouh}
\address{Laboratoire de Syst\`emes Dynamiques et G\'eom\'etrie, Facult\'e de Math\'ematiques, Universit\'e de Sciences et Technologie Houari Boumedienne, Bab Zouar, Alger, Algeria.}
\email{hbanouh@usthb.dz}
\author[Ben Mabrouk]{Anouar Ben Mabrouk}
\address{Higher institut of Applied Mathematics and Computer Sciences, University of Kairaouan, Street of Assad Ibn Alfourat, Kairaouan 3100, Tunisia.\\
and\\
Laboratory of Algebra, Number Theory and Nonlinear Analysis LR18ES15, Department of Mathematics, Faculty of Sciences, Monastir, Tunisia.\\
and\\
Department of Mathematics, Faculty of Sciences, University of Tabuk, King Faisal Rd, Tabuk, Saudi Arabia.}
\email{anouar.benmabrouk@fsm.rnu.tn}
\subjclass{30G35, 42C40, 42B10, 15A66.}
\keywords{Harmonic analysis, Clifford algebra, Clifford analysis, Continuous wavelet transform, Clifford-Fourier transform, Clifford-wavelet transform, Uncertainty principle.}

\begin{abstract} 
In the present work we are concerned with the development of a new uncertainty principle based on wavelet transform in the Clifford analysis/algebras framework. We precisely derive a sharp Heisenberg-type uncertainty principle for the continuous Clifford wavelet transform. 
\end{abstract}
\maketitle
\section{Introduction}
In the last decade(s) a coming back to Clifford algebras and analysis has taken place in many fields in mathematics, physiscs as well as computer science. This may be due to the fact that Clifford algebras incorporate inside one single structure the geometrical and algebraic properties of Euclidean space. For backgrounds on Clifford algebras/analysis the readers may refer to the original works  \cite{Clifford1882,Dirac1928,Grossmann1986,Grossmann1984,Hamilton1844,Hamilton1866,Weyl1950} which are already free.

Insite of a Clifford algebra geometric entities are treated ccording to their dimension such as scalars, vectors, bivectors and volume elements. This has allowed researchers especially in mathematical physics to develop harmonic analysis on Clifford algebras as extension of complex one.  Famous operators and related inequalities such as Cauchy-Riemann operator and uncertainty principle have been extended to the Clifford framework. 

Recently and especially with the research group of Ghent university (Sommen, Brackx, Delange, De Schepper, Hitzer, Mawardi, and their collaborators) have extended wavelet theory to the Clifford framework. Next, a series of works have been developed in the same topic. See \cite{Arfaoui-Rezgui-BenMabrouk,Arfaoui-BenMabrouk1,Arfaoui-BenMabrouk2,Arfaoui-BenMabrouk-Cattani1,Arfaoui-BenMabrouk-Cattani2,BrackxChisholmSoucek,Brackx-Schepper-Sommen0,Brackxetal2006,Brackx-Schepper-Sommen2,Brackxetal2006a,Brackxetal2013,Brackx2001a,Bujacketal,Bujacketal1,Bujacketal2,Delanghe,DeBie1,DeBie2,DeSchepper2006,Hitzer,Hitzer1,Hitzer2,Hitzer-Mawardi-1,Hitzer-Tachibana,Mawardi,Mawardi-Hitzer-1,Mawardi-Ashino-Vaillancourt,Sommen2015,Wietzke,Zou}. 

In the original theory of wavelets on $\mathbb{R}$ as well as its extension to other more general contexts, wavelet analysis of functions starts by checking that the mother wavelet is admissible or not to guarantee the most useful rules by the next such as the reconstruction, Fourier-Plancherel and Parseval ones. For real wavelet analysis, this is explained as follows. Given an analyzing function $\psi\in L^2(\mathbb{R})$ which will be called by the next the mother wavelet, such a function $\psi$ has to satisfy the admissibility condition  
\[
\mathcal{A}_{\psi}=\int_{\mathbb{R}}\frac{\left|\widehat{\psi}(\xi)\right|^{2}}{\left|\xi\right|}d\xi<+\infty,
\]
where $\widehat{\psi}$ is the classical Fourier transform of $\psi$.

The present work is concerned with the development of a sharp uncertainty principle in the Clifford framework based on Clifford wavelets and thus extend the one developed by Banouh et al in \cite{Banouh}. 

As we have mentioned above, the connection between wavelet theory and Clifford algebra/analysis is now growing up and applications are widespread such as in signal/image processing. See \cite{Brackx2001a,DeSchepper2006,Sommen2015,Rizo}. Applications also in physics, nuclear magnetics, electric engineering, color images may be found in \cite{Dian,Hitzer,Rizo,Sauetal,Soulard,Wietzke}.

The aim of the present paper is to establish a new Heisenberg uncertainty principle
applied to continuous Clifford wavelet transform in the settings of the non commutative Clifford algebras. In section 2 some preliminaries on Clifford toolkits and wavelets are presented. Section 3 is concerned to a review on the uncertainty principle in both its classical forms and Clifford one. Section 4 is devoted to the development of our main results. 
\section{Clifford wavelets toolkit}
In this section, we review briefly the basic concepts of Clifford analysis and the continuous wavelet transform on the real Clifford algebra. 
\subsection{Clifford analysis/algebra toolkit}
The Clifford algebra $\mathbb{R}_{n}$ associated to $\mathbb{R}^{n}$ is the $2^n$-dimensional non commutative associative algebra generated by the canonical basis $\left\{e_{1},e_{2},\dots,e_{n}\right\} $ relatively to the non commutative product
\[
e_{i}e_{j}+e_{j}e_{i}=-2\delta_{ij},
\]
where $\delta$ is the Kronecker symbol. Any element $a\in\mathbb{R}_{n}$ has a unique representation 
\[
a=\sum_{A}a_{A}e_{A},\;a_{A}\in\mathbb{R}=\sum_{k=0}^{n}\sum_{\left|A\right|=k}a_{A}e_{A},
\]
where $|A|$ is the length of the multi-index $A$.\\
On the algebra $\mathbb{R}_{n}$ there are some types of involutive operators defined on the basis elements. A main-involution satisfying $\widetilde{e_{A}}=(-1)^{\left|A\right|}e_{A}$. A second type called reversion stating that $e_{A}^{\ast}=(-1)^{\frac{\left|A\right|(\left|A\right|-1)}{2}}e_{A}$ and finally a conjugation operator defined by $\overline{e_{A}}=(-1)^{\frac{|A|(|A|+1)}{2}}e_{A}$. In the case of the complex extension $\mathbb{C}_{n}=\mathbb{R}_{n}+i\mathbb{R}_{n}$ the last operator has to be extended to the so-called Hermitian conjugation defined by $\lambda^{\dagger}=\overline{a}-i\overline{b}$ for $\lambda=a+ib\in\mathbb{C}_{n}$ with $a,b\in\mathbb{R}_{n}$. 

As the vextor space $\mathbb{R}^{n}$ is now seen as a subspace of $\mathbb{R}_{n}$ any element $x=(x_{1},x_{2},\dots,x_{n})\in\mathbb{R}^{n}$
is also seen as an element in $\mathbb{R}_{n}$. We precisely identify it to $\underline{x}={\displaystyle \sum_{j=1}^{n}x_{j}e_{j}}$. This allows to introduce the Clifford product of vectors by
$$
\underline{x}\underline{y}=\underline{x}\cdot\underline{y}+\underline{x}\wedge\underline{y},
$$
where 
$$
\underline{x}\cdot\underline{y}=-<\underline{x},\underline{y}>=-{\displaystyle \sum_{j=1}^{n}x_{j}y_{j}},
$$
and 
$$
\underline{x}\wedge\underline{y}={\displaystyle\sum_{j<k}e_{j}e_{k}(x_{j}y_{k}-x_{k}y_{j})}.
$$
This yields particularly that 
$$
\underline{x}^{2}=-\left|\underline{x}\right|^{2}=-\sum_{j=1}^{n}|x_{j}|^{2}.
$$
Finally we recall that for two functions $f,g:\mathbb{R}^n\longrightarrow\mathbb{C}_n$ the inner product is defined by
\begin{equation}\label{eq:inner product}
<f,g>_{L^{2}(\mathbb{R}_{n},dV(\underline{x}))} =\int_{\mathbb{R}^{n}}\left[f(\underline{x})\right]^{\dagger}g(\underline{x})dV(\underline{x})
\end{equation}
and the associated norm by
$$
\|f\|_{L^{2}(\mathbb{R}_{n},dV(\underline{x}))} =<f,f>_{L^{2}(\mathbb{R}_{n},dV(\underline{x}))}^{\frac{1}{2}}.
$$
We denote also
\[
\left|f\right|_{L^{1}(\mathbb{R}_{n},dV(\underline{x}))}=\int_{\mathbb{R}^{n}}|f(\underline{x})|dV(\underline{x})
\]
where $dV(\underline{x})$ stands for the Lebesgue measure. As in the classical cases, it is easy to show the Cauchy-Schwartz inequality
\begin{equation}
\left|<f,g>_{L^{2}(\mathbb{R}_{n},dV(\underline{x}))}\right|\leq\left|f\right|_{L^{2}(\mathbb{R}_{n},dV(\underline{x}))}\left|g\right|_{L^{2}(\mathbb{R}_{n},dV(\underline{x}))}.\label{cauchy schwartz}
\end{equation}
\subsection{The continuous Clifford-wavelet transform}
As for multidimensional Euclidean spaces the Fourier and wavelet transforms have been extended to the case of Clifford framework. The Clifford-Fourier transform of a Clifford-valued function $f\in L^{1}\cap L^{2}(\mathbb{R}_{n},dV(\underline{x}))$
is defined by
\[
\mathcal{F}\left[f\right](\underline{\xi})=\widehat{f}(\underline{\xi})=\frac{1}{(2\pi)^{\frac{n}{2}}}\int_{\mathbb{R}^{n}}e^{-i<\underline{x},\underline{\xi}>}f(\underline{x})dV(\underline{x}).
\]
Here also we may speak about its inverse defined by 
\[
\mathcal{F}^{-1}\left[f\right](\underline{x})=\frac{1}{(2\pi)^{\frac{n}{2}}}\int_{\mathbb{R}^{n}}e^{i<\underline{x},\underline{\xi}>}\widehat{f}(\underline{\xi})dV(\underline{\xi}).
\]
The concept of Clifford-wavelet transform has been now widely studied. In the present paper we will recall a special case based on the spin theory which will be used in our work and which has been already applied in \cite{Banouh}. Let $\psi\in L^{1}\cap L^{2}(\mathbb{R}^{n},dV(\underline{x}))$ be such that 
\begin{itemize}
\item ${\psi}$ is a Clifford-algebra-valued function.
\item $\widehat{\psi}(\underline{\xi})\left[\widehat{\psi}(\underline{\xi})\right]^{\dagger}$ is scalar.
\item $\mathcal{A}_{\psi}=\displaystyle(2\pi)^{n}\int_{\mathbb{R}^{n}}\frac{\widehat{\psi}(\underline{\xi})\left[\widehat{\psi}(\underline{\xi})\right]^{\dagger}}{|\underline{\xi}|^{n}}dV(\underline{\xi})<\infty.$
\end{itemize}
$\psi$ is called by the next a Clifford mother wavelet. We also recall in the context of Clifford framework, the spin group of order $n$ which is defined by
\[
Spin(n)=\left\{ s\in\mathbb{R}_{n};\;s={\displaystyle \prod\limits _{j=1}^{2l}}\underline{\omega}_{j},\;\underline{\omega}_{j}^{2}=-1,1\leq j\leq2l\right\}.
\]
For $(a,\underline{b},s)\in\mathbb{R}_{+}\times\mathbb{R}^{n}\times Spin(n)$,
we denote
\[
\psi^{a,\underline{b},s}(\underline{x})=\frac{1}{a^{\frac{n}{2}}}s\psi(\frac{\overline{s}(\underline{x}-\underline{b})s}{a})\overline{s}.
\]
In \cite{Banouh} the quthors showed that the copies ${\psi^{a,\underline{b},s}}$ and that they constitute a dense set in $L^{2}(\mathbb{R}^{n},dV(\underline{x}))$.
\begin{defn} The Clifford-wavelet transform of an analyzed function $f\in L^2(\mathbb{R}^n,dV(\underline{x}))$ is defined by
	\[
	T_{\psi}\left[f\right](a,\underline{b},s)=\int_{\mathbb{R}^{n}}\left[\psi^{a,\underline{b},s}(\underline{x})\right]^{\dagger}f(\underline{x})dV(\underline{x}).
	\]
Let $\mathcal{H}_{\psi}=T_{\psi}\left(L^{2}(\mathbb{R}^{n},dV(\underline{x}))\right)$. We define an inner product on $\mathcal{H}_{\psi}$ by
\[
\left[T_{\psi}\left[f\right],T_{\psi}\left[g\right]\right]=\frac{1}{\mathcal{A}_{\psi}}\int\limits_{Spin(n)}\int\limits_{\mathbb{R}^{n}}\int\limits _{\mathbb{R}_{+}}(T_{\psi}\left[f\right](a,\underline{b},s))^{\dagger}T_{\psi}\left[g\right](a,\underline{b},s)d\mu(a,\underline{b},s),
\]
where $d\mu(a,\underline{b},s)=\frac{da}{a^{n+1}}dV(\underline{b})ds$ and where $ds$ is the Haar measure on $Spin(n)$.
\end{defn}
It is straightforwrd that the operator $T_{\psi}$ is an isometry from  $L^{2}(\mathbb{R}^{n},dV(\underline{x}))$ to $\mathcal{H}_{\psi}$. See \cite{Banouh}. As a result any analyzed function $f\in L^{2}(\mathbb{R}^{n},dV(\underline{x}))$ may be reconstructed in the $L^2$-sense as 
\[
f(\underline{x})=\frac{1}{\mathcal{A}_{\psi}}\int\limits _{Spin(n)}\int\limits _{\mathbb{R}^{n}}\int\limits _{\mathbb{R}_{+}}\psi^{a,\underline{b},s}(\underline{x})T_{\psi}\left[f\right](a,\underline{b},s)\frac{da}{a^{n+1}}dV(\underline{b})ds.
\]
This constitutes the well-known Plancheral-Parsevall formula extended to the Clifford wavelet framework. 
\subsection{The uncertainty principle revisited}
The uncertainty principle is originally due to Heisenberg (See \cite{Heisenberg1927}, \cite{Heisenberg1985}) and constitutes since its discovery an interesting concepts in quantum mechanics. Mathematically, it is resumed by the fact that a non-zero function and its Fourier transform cannot both be sharply localized. 

In the last decades especially of the last century many studies have been developed tackling the uncertainty principle. Many variants have been thus proposed using pure Fourier transform, pure wavelets, Clifford Fourier transform and recently Clifford wavelet transform. In \cite{Amri-Rachdi-1} for example a Stein–Weiss type uncertainty inequality has been proved. In \cite{Amri-Rachdi-2} Fourier transform combined with Riemann–Liouville operator has been applied for a Hausdorff–Young inequality to yield an entropy based uncertainty principle and a Heisenberg–Pauli–Weyl inequality. In \cite{Rachdi-Meherzi} continuous wavelet transform has been applied to prove an analogue of Heisenberg's inequality (See also \cite{Rachdi-Amri-Hammami}, \cite{Rachdi-Herch}).

In \cite{Feichtinger-Grochenig} using Heisenberg group techniques, stable iterative algorithms for signal analysis and synthesis the authors established a variant of the Heisenberg uncertainty inequality. In \cite{Dahkleteal} an uncertainty principle has been shown using shearlet transform. El-Haoui et al in \cite{ElHaouietal} applied a quaternionic Fourier transform to derive an uncertainty principles including Heisenberg-Weyls, Hardys, Beurlings and logarithmic ones. See also \cite{ElHaoui-Fahlaoui}.

Next, in the Clifford framework Hitzer and collaborators were the most active group in developing uncertainty principle variants such as \cite{Hitzer1,Hitzer2} where some generalized Clifford wavelet uncertainty principles have been proposed. See also \cite{Hitzer-Tachibana}. Hitzer and Mawardi in \cite{Hitzer-Mawardi-1} extended the Fourier transform to some general Clifford geometric algebras leading to a a new uncertainty principle for multivector functions in the new Clifford geometric algebras. See also \cite{Mawardi-Ryuichi}, \cite{Mawardi-Ryuichi-1}, \cite{Mawardi-Ryuichi-2}, \cite{Mawardi-Hitzer-1}, \cite{Mawardi-Hitzer-2}, \cite{Mawardi-Hitzer-3}, \cite{Mawardi-Hitzer-4}, \cite{Mawardi-Hitzer-Hayashi-Ashino}, \cite{Mawardi-Ashino-Vaillancourt}.

In \cite{Ma-Zhao} the authors used quaternion ridgelet transform and curvelet transform associated to the quaternion Fourier transform to establish an uncertainty principle. Kou et al in \cite{Kouetal} applied the linear quaternion canonical transform has been applied to establish an uncertainty principle. Yang et al in \cite{Yangetal1} have investigated a stronger directional uncertainty principles based on Fourier transform.

For about the uncertainty principle and different variants based on Fourier and wavelet transforms may be found in \cite{Jorgensen}, \cite{Nagata}, \cite{Sen}, \cite{Stabnikov}, \cite{Yang-Kou}.

We propose by the next to recall some mathematical formulations of the uncertainty principle from both classical and recent forms. 
\begin{thm}\label{uncertainty-cas-general} (\cite{Banouh,Weyl1950}) Let $A$ and $B$ be two self-adjoint operators on a Hilbert space $X$ with domains $\mathcal{D}(A)$ and $\mathcal{D}(B)$ respectively and consider their commutator $\left[A,B\right]=AB-BA$. Then
\begin{equation}
\left|Af\right|_{2}\left|Bf\right|_{2}\geq\displaystyle\frac{1}{2}\left|<\left[A,B\right]f,f>\right|,\forall f\in\mathcal{D}(\left[A,B\right]).
\end{equation}
\end{thm}
Consider next the special case 
$$
A_{k}f(\underline{x})=x_{k}f(\underline{x})\;\;\hbox{and}\;\;
B_{k}f(\underline{x})=\partial_{x_{k}}f(\underline{x}),\;k=1,2,\cdots,n.
$$
By applying Theorem \ref{uncertainty-cas-general} above, we get the following result.
\begin{cor}\label{cor1}(\cite{Banouh}) 
\[
\left|A_{k}f\right|_{L^{2}(\mathbb{R}_{n},dV(\underline{x}))}\left|B_{k}f\right|_{L^{2}(\mathbb{R}_{n},dV(\underline{x}))}\geq\frac{1}{2}\left|<\left[A_{k},B_{k}\right]f,f>\right\vert.
\]
Furthermore,
\begin{equation}\label{uncer clifford-fourier}
\|x_{k}f\|_{L^{2}(\mathbb{R}_{n},dV(\underline{x}))}\|\xi_{k}\widehat{f}\|_{L^{2}(\mathbb{R}_{n},dV(\underline{x}))}\geq\frac{1}{2}\|f\|_{L^{2}(\mathbb{R}_{n},dV(\underline{x}))}^{2}.
\end{equation}
\end{cor}
Moreover, by considering the oerator $T_{\psi}$ the authors is \cite{Banouh} established the follozing result.
\begin{thm}\label{main-result-Banouh}(\cite{Banouh}) Let be $\psi\in L^{2}(\mathbb{R}^{n},\mathbb{R}_{n},dV(\underline{x}))$ be an admissible Clifford mother wavelet. Then for $f\in L^{2}(\mathbb{R}^{n},\mathbb{R}_{n},dV(\underline{x}))$. It holds for $1\leq k\leq n$ that 
\begin{equation}
{\displaystyle \left({\displaystyle \int_{Spin(n)}{\displaystyle \int_{\mathbb{R}^{+}}\left|b_{k}T\left[f\right](a,\cdot,s)\right|_{2}^{2}\frac{da}{a^{n+1}}ds}}\right)^{\frac{1}{2}}\left|\xi_{k}\widehat{f}\right|_{2}\geq\frac{(2\pi)^{\frac{n}{2}}}{2}\sqrt{A_{\psi}}\left|f\right|_{2}^{2}.}
\end{equation}
\end{thm} 
\section{A sharper Clifford wavelet uncertainty principle}
In the present section we state and prove our main result in the present work. It consists a sharper formulation of the Clifford-wavelet uncertainty principle. The main result is subject of the following theorem.
\begin{thm}\label{new-main-result}
Let $\psi\in L^{2}(\mathbb{R}^{n},\mathbb{R}_{n},dV(\underline{x}))$ be an admissible Clifford mother wavelet. Then for $f\in L^{2}(\mathbb{R}^{n},\mathbb{R}_{n},dV(\underline{x}))$, the following uncertainty principle holds,
\[
\begin{array}{ccc}
\left(\displaystyle\int_{Spin(n)}\int_{\mathbb{R}^{+}}\left\|b_{k}T_{\psi}\left[f\right](a,\cdot,s)\right\|^{2}\frac{da}{a^{n+1}}ds\right)^{\frac{1}{2}}\left\|\xi_{k}\widehat{f}\right\Vert\\
\\
\qquad\qquad\qquad\qquad\qquad\qquad\qquad \geq\sqrt{2^{n+1}\pi{}^{n}A_{\psi}}\left\{\left\|f\right\|^{2}+2\left|\left\langle f_1,f_2\right\rangle \right|\right\}
\end{array}
\]
where
$$
f_1(\underline{x})=\frac{1}{A_{\psi}}\int_{Spin(n)}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{+}}\psi^{a,\underline{b},s}(\underline{x})\partial_{b_{k}}T_{\psi}\left[f\right](a,\underline{b},s)\frac{da}{a^{n+1}}dV(\underline{b})ds
$$
and
$$
f_2(\underline{x})=\frac{1}{A_{\psi}}\int_{Spin(n)}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{+}}\psi^{a,\underline{b},s}(\underline{x})b_{k}T_{\psi}\left[f\right](a,\underline{b},s)\frac{da}{a^{n+1}}dV(\underline{b})ds.
$$
\end{thm}
\noindent\textbf{Proof.} Observe firstly that 
\[
\left\|x_{k}f\right\|\left\|\xi_{k}\widehat{f}\right\|\geq\sqrt{2}(\left\|f\right\|^{2}+\left|2\left\langle x_{k}\partial_{x_{k}}f,f\right\rangle \right|).
\]
We substitute $f(\bullet)$ by $T_{\psi}\left[f\right](a,\bullet,s)$. We obtain
$$
\begin{array}{lll}
\left\|b_{k}T_{\psi}\left[f\right](a,\bullet,s)\right\|\left\|\xi_{k}\widehat{T_{\psi}\left[f\right](a,\bullet,s)}\right\|\\ \qquad\qquad\qquad\qquad\qquad\geq\sqrt{2}(\left\|T_{\psi}\left[f\right](a,\bullet,s)\right\|^{2}+\mathcal{S}_{\psi}\left[f\right](a,\bullet,s))
\end{array}
$$
where 
$$
\mathcal{S}_{\psi}\left[f\right](a,\bullet,s)=\left|2\left\langle b_{k}\partial_{b_{k}}T_{\psi}\left[f\right](a,\bullet,s),T_{\psi}\left[f\right](a,\bullet,s)\right\rangle \right|.
$$
As a result,
$$
\begin{array}{lll}
&\displaystyle\int_{Spin(n)}\int_{\mathbb{R}^{+}}\left\|b_{k}T_{\psi}\left[f\right](a,\bullet,s)\right\|\left\|\xi_{k}\widehat{T_{\psi}\left[f\right](a,\bullet,s)}\right\|\frac{da}{a^{n+1}}ds\\
\\
&\quad\geq\sqrt{2}\displaystyle\int_{Spin(n)}\int_{\mathbb{R}^{+}}(\left\|T_{\psi}\left[f\right](a,\bullet,s)\right\|^{2}+\mathcal{S}_{\psi}\left[f\right](a,\bullet,s))\dfrac{da}{a^{n+1}}ds.
\end{array}
$$
By Cauchy-Schwartz we get
$$
\begin{array}{lll}
&\displaystyle\int_{Spin(n)}\int_{\mathbb{R}^{+}}\left\|b_{k}T_{\psi}\left[f\right](a,\bullet,s)\right\|\left\Vert \xi_{k}\widehat{T_{\psi}\left[f\right](a,\bullet,s)}\right\Vert \frac{da}{a^{n+1}}ds\\
\\
&\qquad\qquad\displaystyle\leq(\int_{Spin(n)}\int_{\mathbb{R}^{+}}\left\Vert b_{k}T_{\psi}\left[f\right](a,\bullet,s)\right\Vert ^{2}\dfrac{da}{a^{n+1}}ds)^{\frac{1}{2}}\\
\\
&\qquad\qquad\qquad\displaystyle\times(\int_{Spin(n)}\int_{\mathbb{R}^{+}}\left\Vert \xi_{k}\widehat{T_{\psi}\left[f\right](a,\bullet,s)}\right\Vert ^{2}\dfrac{da}{a^{n+1}}ds)^{\frac{1}{2}}
\end{array}
$$
Denote next 
$$
BT_{\psi}\left[f\right]=(\int_{Spin(n)}\int_{\mathbb{R}^{+}}\left\Vert b_{k}T_{\psi}\left[f\right](a,\bullet,s)\right\Vert ^{2}\frac{da}{a^{n+1}}ds)^{\frac{1}{2}}
$$
and
$$
\xi T_{\psi}\left[f\right]=(\int_{Spin(n)}\int_{\mathbb{R}^{+}}\left\Vert \xi_{k}\widehat{T_{\psi}\left[f\right](a,\bullet,s)}\right\Vert ^{2}\frac{da}{a^{n+1}}ds)^{\frac{1}{2}}.
$$
Then, we get 
\begin{align*}
BT_{\psi}\left[f\right]\times\xi T_{\psi}\left[f\right]
&\geq\sqrt{2}\int_{Spin(n)}\int_{\mathbb{R}^{+}}(\left\Vert T_{\psi}\left[f\right](a,\bullet,s)\right\Vert ^{2}\\
&\quad\quad+\left|2\left\langle b_{k}\partial_{b_{k}}T_{\psi}\left[f\right](a,\bullet,s),T_{\psi}\left[f\right](a,\bullet,s)\right\rangle \right|)\frac{da}{a^{n+1}}ds\\
&=\sqrt{2}\int_{Spin(n)}\int_{\mathbb{R}^{+}}(\left\Vert T_{\psi}\left[f\right](a,\bullet,s)\right\Vert ^{2}\frac{da}{a^{n+1}}ds\\
&\quad\quad+\sqrt{2}\int_{Spin(n)}\int_{\mathbb{R}^{+}}\mathcal{S}_{\psi}\left[f\right](a,\bullet,s)\frac{da}{a^{n+1}}ds.
\end{align*}
Observe next that 
$$
\int_{Spin(n)}\int_{\mathbb{R}^{+}}\left\Vert \xi_{k}\widehat{T_{\psi}\left[f\right]}(a,\bullet,s)\right\Vert ^{2}\frac{da}{a^{n+1}}ds=\frac{A_{\psi}}{(2\pi)^{n}}\left\Vert \xi_{k}\widehat{f}\right\Vert ^{2}
$$
and that 
$$
\int_{Spin(n)}\int_{\mathbb{R}^{+}}\left\Vert T_{\psi}\left[f\right](a,\bullet,s)\right\Vert ^{2}\frac{da}{a^{n+1}}ds =A_{\psi}\left\Vert f\right\Vert ^{2}
$$
and denote 
$$
\widetilde{BT_{\psi}}\left[f\right]=
\int_{Spin(n)}\int_{\mathbb{R}^{+}}\left|\left\langle b_{k}\partial_{b_{k}}T_{\psi}\left[f\right](a,\bullet,s),T_{\psi}\left[f\right](a,\bullet,s)\right\rangle \right|)\frac{da}{a^{n+1}}ds.
$$
We deduce that  
\begin{align*}
BT_{\psi}\left[f\right]\times(\frac{A_{\psi}}{(2\pi)^{n}}\left\Vert \xi_{k}\widehat{f}\right\Vert ^{2})^{\frac{1}{2}}\geq\sqrt{2}A_{\psi}\left\Vert f\right\Vert ^{2}+2\sqrt{2}\widetilde{BT_{\psi}}\left[f\right].
\end{align*}
Consequently,
$$
\begin{array}{lll}
&\displaystyle(\int_{Spin(n)}\int_{\mathbb{R}^{+}}\left\Vert b_{k}T_{\psi}\left[f\right](a,\bullet,s)\right\Vert ^{2}\frac{da}{a^{n+1}}ds)^{\frac{1}{2}}\left\Vert \xi_{k}\widehat{f}\right\Vert\\
\\
&\displaystyle\geq\sqrt{2^{n+1}\pi{}^{n}A_{\psi}}\left\Vert f\right\Vert^{2}
+\displaystyle\sqrt{\frac{2^{n+1}\pi{}^{n}}{A_{\psi}}}\int_{Spin(n)}\int_{\mathbb{R}^{+}}\mathcal{S}_{\psi}\left[f\right](a,\bullet,s)\dfrac{da}{a^{n+1}}ds.
\end{array}
$$
Now, observe that
$$
\begin{array}{lll}
\displaystyle\int_{Spin(n)}\int_{\mathbb{R}^{+}}\left|\left\langle b_{k}\partial_{b_{k}}T_{\psi}\left[f\right](a,\bullet,s),T_{\psi}\left[f\right](a,\bullet,s)\right\rangle \right|)\frac{da}{a^{n+1}}ds\\
\\
=\displaystyle\int_{Spin(n)}\int_{\mathbb{R}^{+}}\left|\int_{\mathbb{R}^{n}}\left[\partial_{b_{k}}T_{\psi}\left[f\right](a,\underline{b},s)\right]^{\dagger}b_{k}T_{\psi}\left[f\right](a,\underline{b},s)dV(\underline{b})\right|\frac{da}{a^{n+1}}ds\\
\\
\geq\displaystyle\left|\int_{Spin(n)}\int_{\mathbb{R}^{+}}\int_{\mathbb{R}^{n}}\left[\partial_{b_{k}}T_{\psi}\left[f\right](a,\underline{b},s)\right]^{\dagger}b_{k}T_{\psi}\left[f\right](a,\underline{b},s)dV(\underline{b})\frac{da}{a^{n+1}}ds\right|\\
\\
=\displaystyle\left|A_{\psi}\left[\partial_{b_{k}}T_{\psi}\left[f\right],b_{k}T_{\psi}\left[f\right]\right]\right|.
\end{array}
$$
We obtain  
$$
\begin{array}{lll}
\displaystyle(\int_{Spin(n)}\int_{\mathbb{R}^{+}}\left\Vert b_{k}T_{\psi}\left[f\right](a,\bullet,s)\right\Vert ^{2}\frac{da}{a^{n+1}}ds)^{\frac{1}{2}}\left\Vert \xi_{k}\widehat{f}\right\Vert\\
\\
\qquad\qquad\geq\sqrt{2^{n+1}\pi{}^{n}A_{\psi}}\left\Vert f\right\Vert ^{2}
+\sqrt{2^{n+3}\pi{}^{n}A_{\psi}}\left|\left[\partial_{b_{k}}T_{\psi}\left[f\right],b_{k}T_{\psi}\left[f\right]\right]\right|.
\end{array}
$$
Since for $f,g\in L^{2}(\mathbb{R}^{n},dV(\underline{x}))$ it holds that
\[
\left[T_{\psi}\left[f\right],T_{\psi}\left[g\right]\right]=\left\langle f,g\right\rangle 
\]
we may write 
$$
\begin{array}{lll}
\displaystyle(\int_{Spin(n)}\int_{\mathbb{R}^{+}}\left\Vert b_{k}T_{\psi}\left[f\right](a,\bullet,s)\right\Vert ^{2}\frac{da}{a^{n+1}}ds)^{\frac{1}{2}}\left\Vert\xi_{k}\widehat{f}\right\Vert\\
\\
\qquad\qquad\qquad\qquad\geq\sqrt{2^{n+1}\pi{}^{n}A_{\psi}}\left\{\left\Vert f\right\Vert ^{2}+2\left|\left\langle f_1,f_2\right\rangle \right|\right\}
\end{array}
$$
where
$$
f_1(\underline{x})=\frac{1}{A_{\psi}}\int_{Spin(n)}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{+}}\psi^{a,\underline{b},s}(\underline{x})\partial_{b_{k}}T_{\psi}\left[f\right](a,\underline{b},s)\frac{da}{a^{n+1}}dV(\underline{b})ds
$$
and
$$
f_2(\underline{x})=\frac{1}{A_{\psi}}\int_{Spin(n)}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{+}}\psi^{a,\underline{b},s}(\underline{x})b_{k}T_{\psi}\left[f\right](a,\underline{b},s)\frac{da}{a^{n+1}}dV(\underline{b})ds.
$$
\section{Conclusion}
In this paper, an uncertainty principle based on the continuous wavelet transform in the Clifford algebra's settings has been formulated and proved. Compared to xisting formulations in Clifford framework, the proposed uncertqinty principle here is sharper.
\section{Data availability statement}
Data sharing is not applicable to this article as no new data were created or analyzed in this study.

\end{document}